solve 3x+3y+3z=0, -3x-3y-3z=0 and 3x+3y+3z=o
i have tried it but i am not getting the expected answer i,e. -1 1 0 please can you tell me the way to solve this

Answer by Alan3354(69443)

3x+3y+3z=0, -3x-3y-3z=0 and 3x+3y+3z=o
------------
That's the same equation 3 times.
No unique solution.
(-1,-1,0) works, but so does (1,1,-2) and (2,4,-6)